OpenStudy (anonymous):

Rachel and Sarah belong to different gyms. Rachel's gym charges a one-time registration fee of $15 and then charges $35 per month. Whereas Sarah only pays $25 per month but had to pay a $75 registration fee. After how many months will Rachel and Sarah have spent the same amount on their gym memberships?

OpenStudy (marigirl):

Develop two equations, then you can either graph it and find its intercepts or make both the equations equal to each other and solve for x Rachel: y=35x+15 where x stands for the number of months Sarah: y=25x+75

OpenStudy (marigirl):

Then we can make both equations equal to each other to find the point of intersection (where both gyms would cost the same) \[35x+15=25x+75\] \[35x-25x=75-15\] \[10x=60\] \[x=6\] Therefore it both the gyms would cost the same if they wish to attend for up to 6 months.

OpenStudy (marigirl):